# The bispectral problem, the Darboux process, monodromy and the Hermite operator

### v1

@@ 1. A brief introduction
The results in [19] feature two different situations: a pair of “rank one” bispectral families and a “rank two” bispectral family. Both of these are obtained by repeated applications of what is known as the Darboux process to a second order differential operator of the form
@@
The Darboux process can be applied to difference operators as well, as first done in [49], and its usefulness in the context of the bispectral problem was exploited in [37, 38, 39, 40, 41]. Of particular interest in these papers is the case of the Hermite polynomials. This is the one case, among the classical orthogonal polynomials, where the support of the orthogonality measure runs over the entire real line making it quite different from the case of the Krall-Jacobi and the Krall-Laguerre polynomials discussed in the first three papers just cited. The application of the Darboux process is greatly simplified when the orthogonality measure has at least one finite end point, and this feature sets the Hermite case apart from the other ones. This is treated in detail in [40]. See also [18] for a way to handle a related issue.
We close this introduction by noticing that the bispectral property in [19] was initially considered as a step in an ambitious program dealing with the issue of “time-and-band-limiting” as motivated by medical imaging in the presence of limited and noisy data. The introduction to [19] briefly mentions this; for further work one can see [33, 34, 35] as well as the more recent set of papers [7, 8, 9, 10] together with [11, 12, 13, 14, 15, 42, 43, 44].
@@ 2. The contents of the paper
One more comment on [19]. The big surprise in this paper is the connection between the bispectral problem and integrable systems such as KdV. Many of the tools in this paper were “floating around” in connection with developments around the KdV equation around the 1970s, and this allowed the authors to use these tools. There are two tools that appear for the first time in this paper: the relevance of the ad-conditions and of monodromy considerations in connection with the bispectral property, in particular the characterization given in Proposition 3.3 and its consequence Theorem 3.4, page 196. One might imagine that this Proposition is a result from the 19th century, however see [23, 51, 56]. Interesting extensions to the matrix valued case have been given by A. Veselov and his collaborators. This result has been used, in the scalar valued case, for instance in [32, 51].
@@ 3. The Hermite operator
$x\ _{1}F_{1}\left(\frac{1}{2}-\frac{n}{2};\frac{3}{2};x^{2}\right)=\ _{1}F_{1}\left(-\frac{n}{2};\frac{1}{2};x^{2}\right)\int_{0}^{x}\dfrac{e^{y^{2}}}{{}_{1}F^{2}_{1}\left(-\frac{n}{2};\frac{1}{2};y^{2}\right)}dy$
@@
$\xi_{n}(\pm ix)=\dfrac{1}{\sqrt{\pi}n!}\left[\Gamma\left(\frac{n+1}{2}\right)\,_{1}F_{1}\left(-\frac{n}{2};\frac{1}{2};x^{2}\right)\mp 2ix\Gamma\left(\frac{n}{2}+1\right)\,_{1}F_{1}\left(\frac{1}{2}-\frac{n}{2};\frac{3}{2};x^{2}\right)\right]$
@@
(7) $\psi_{2n+1}(x)$ $=$ $(-1)^{n+1}4^{n}n!\,_{1}F_{1}\left(-n-\frac{1}{2};\frac{1}{2};x^{2}\right),$
(8) $\psi_{2n}(x)$ $=$ $(-4)^{n}n!x\,_{1}F_{1}\left(\frac{1}{2}-n;\frac{3}{2};x^{2}\right),\quad n\geq 0.$
@@
The results in [19] cannot be taken as a general rule in terms of the relation between the Darboux process going from one differential operator to another and the corresponding monodromy groups. The peculiar situation in [19] is that a sequence of Darboux transformations preserves the monodromy group in two different instances, in one case the groups are all trivial and in the other case the coincide with the integers. The paper [26] looks at an example of the hypergeometric equation of Euler, Gauss and Kummer with a finite monodromy group (all of these were characterized by H. Schwartz) and one sees that an application of the Darboux process to it gives a differential operator with an infinite monodromy group.
@@
This is much simpler than the situation in the Hermite case, there is no need to do any scaling, as in (7) and (8) above starting from the most familiar solutions of the differential equation, namely $P_{n}$ and $Q_{n}$.
@@ 4. Example 1: Hermite with one gap
(10) $\mathcal{H}_{x}+2I=\partial_{x}^{2}-2x\partial_{x}+2=(\dfrac{1}{x}\partial_{x}-2)(x\partial_{x}-1)$
@@
$p_{n}(x)=\det\begin{pmatrix}x&1\\ H_{n}&H^{\prime}_{n}\end{pmatrix}$.
@@
(11) $\dfrac{n(n-1)}{2}p_{n-2}(x)+\dfrac{2n-1}{4}p_{n}(x)+\dfrac{n-1}{8(n+1)}p_{n+2}(x)=\dfrac{x^{2}}{2}p_{n}(x),\quad n\geq 1,$
@@
For a small sample of references on this topic see [22, 24, 28, 30, 31, 48, 52, 55]. It may be worth remarking that the first use of the Darboux process in the context of exceptional orthogonal polynomials appears to be [52] where the author very graciously says “This procedure whose idea goes back to an older paper [19]”, while the first observation of the presence of recursion relations appears to be [55].
@@
highlighting the important role of the operator $x\partial_{x}-1$ that appears in (10) above, in giving new eigenfunctions in terms of old ones.
@@
$\phi_{6}(x)$ $=$ $-120\widetilde{\alpha}_{0}\left(1+6x^{2}-12x^{4}+\dfrac{8}{3}x^{6}\right)+480\widetilde{\beta}_{2}\left(x^{3}-\dfrac{3}{5}x^{5}+\dfrac{3}{70}x^{7}+\ldots\right)$
@@ 5. Example 2: Hermite with two gaps
$q_{n}(x)=\det\begin{pmatrix}4x^{2}-2&8x\\ H_{n-1}&H^{\prime}_{n-1}\end{pmatrix}$.
@@
$\phi_{1}(x)$ $=$ $\widetilde{\alpha}_{1}\left(1+3x^{2}-\dfrac{x^{4}}{6}-\dfrac{13}{30}x^{6}-\dfrac{11}{56}x^{8}+\ldots\right)+\widetilde{\beta_{1}}x$
@@ 6. The ad-conditions
(16) $\left(\textrm{ad}\ X\right)^{n}(Y)=\textrm{ad}\ X\left(\left(\textrm{ad}\ X\right)^{n-1}(Y)\right),\quad n\geq 2.$
@@
$\textrm{ad}\ \mathcal{L}^{3}(\Theta)-4\textrm{ad}\ \mathcal{L}(\Theta)=0,$
@@
$\textrm{ad}\ \mathcal{L}^{4}(\Theta)-16\textrm{ad}\ \mathcal{L}^{2}(\Theta)=0.$

### v2

# arXiv:2509.04158v2[math.CA] 28 Feb 2026

@@ 1. A brief introduction
The results in [19] feature two different situations: a pair of “rank one” bispectral families and a “rank two” bispectral family (see Section 3 for a formal definition of rank in this framework). Both of these are obtained by repeated applications of what is known as the Darboux process to a second order differential operator of the form
@@
The Darboux process can be applied to difference operators as well, as first done in [50], and its usefulness in the context of the bispectral problem was exploited in [36, 37, 38, 39, 40, 41]. Of particular interest in these papers is the case of the Hermite polynomials. This is the one case, among the classical orthogonal polynomials, where the support of the orthogonality measure runs over the entire real line making it quite different from the case of the Krall-Jacobi and the Krall-Laguerre polynomials discussed in the first three papers just cited. The application of the Darboux process is greatly simplified when the orthogonality measure has at least one finite end point, and this feature sets the Hermite case apart from the other ones. This is treated in detail in [40]. See also [18] for a way to handle a related issue.
We notice that the bispectral property in [19] was initially considered as a step in an ambitious program dealing with the issue of “time-and-band-limiting” as motivated by medical imaging in the presence of limited and noisy data. The introduction to [19] briefly mentions this; for further work one can see [32, 33, 35] as well as the more recent set of papers [7, 8, 9, 10] together with [11, 12, 13, 14, 15, 42, 43, 44].
We close this introduction by mentioning that the study of the bispectral problem in the “continuous-continuous” case posed and solved in [19], is full of unexpected surprises that could not be surmised from its simple formulation. Here is a short list:
  • a.
    The ad-conditions are necessary and sufficient for bispectrality.
  • b.
    As a consequence, the potential needs to be rational, one can exclude irregular singularities and the potential needs to decay at infinity. All of this is proved in [19].
  • c.
    The connection with integrable systems (KdV)
  • d.
    The relevance of the Darboux process.
  • e.
    (Most important for the present paper) The relation with trivial or integer monodromy, and the strict connection with rank one and rank two bundles respectively.
In particular the observation about trivial monodromy and the explicit form of the potential gave origin to a paper by V. Goncharenko and A.P. Veselov [26], where they deal with the matrix valued case of the Schroedinger equation and where they say that this is a result that one could imagine was settled in the previous century but was only done in [19]. Reference [52] by A. Oblomkov extends this result from [19] and gives the class of potentials with quadratic growth at infinity that have trivial monodromy. As indicated in the present paper our examples fall within this class.
The work in [19] deals with the continuous-continuous case and sets up a clear correspondence between trivial monodromy and rank one bundles on the one hand and integer monodromy and rank two bundles on the other hand.
The case studied in the present paper is of mixed kind: a potential that explodes at infinity calls for a discrete spectrum and in the best of cases for difference equations satisfied by the eigenfunctions. To the best of our knowledge this situation regarding monodromy has not been explored in the 40 years since the appearence of [19]. What we find here is that the lessons learned earlier do not apply in the case of a very important potential that does not decay at infinity.
Here we display examples with trivial monodromy going with rank two bundles of common eigenfunctions giving a bispectral situation. To the best of our knowledge, this has never been seen before and stands in sharp contrast to the results in [19].
@@ 2. The contents of the paper
One more comment is appropriate here: the big surprise in [19] is the connection between the bispectral problem and integrable systems such as KdV. Many of the tools in this paper were “floating around” in connection with developments around the KdV equation around the 1970s, and this allowed the authors to use these tools. There are two tools that appear for the first time in this paper: the relevance of the ad-conditions and of monodromy considerations in connection with the bispectral property, in particular the characterization given in Proposition 3.3 and its consequence Theorem 3.4, page 196. One might imagine that this Proposition is a result from the 19th century, however see [23, 26, 52]. Interesting extensions to the matrix valued case have been given by A. Veselov and his collaborators. This result has been used, in the scalar valued case, for instance in [31, 52].
@@ 3. The Hermite operator
$x\ _{1}F_{1}\left(\frac{1}{2}-\frac{n}{2};\frac{3}{2};x^{2}\right)=\ _{1}F_{1}\left(-\frac{n}{2};\frac{1}{2};x^{2}\right)\int_{0}^{x}\dfrac{e^{y^{2}}}{{}_{1}F^{2}_{1}\left(-\frac{n}{2};\frac{1}{2};y^{2}\right)}dy.$
@@
$\xi_{n}(\pm ix)=\dfrac{1}{\sqrt{\pi}n!}\left[\Gamma\left(\frac{n+1}{2}\right)\,_{1}F_{1}\left(-\frac{n}{2};\frac{1}{2};x^{2}\right)\mp 2ix\Gamma\left(\frac{n}{2}+1\right)\,_{1}F_{1}\left(\frac{1}{2}-\frac{n}{2};\frac{3}{2};x^{2}\right)\right].$
@@
The rank of the operators in a bispectral situation, such as the one defined by the continuous operator $\mathcal{H}_{x}$ in (4) and the discrete operator $\mathcal{B}_{n}$ in (5), is the dimension of the vector space of common eigenfunctions. In this particular case, for any fixed value of the discrete parameter $n$, $n\geq 1$, the resulting space of solutions in the continuous variable $x$ is two-dimensional, as given by (6). The straightforward definition given above is all that is needed here. However, we mention that this is connected with definitions of rank in the case of commutative rings of differential operators, starting with [6]. A very good account is in [57] (see also [58]) as well as in [46], where Luc Haine reviews joint work of his with Plamen Iliev.
@@
$\psi_{2n+1}(x)$ $=$ $(-1)^{n+1}4^{n}n!\,_{1}F_{1}\left(-n-\frac{1}{2};\frac{1}{2};x^{2}\right),$
$\psi_{2n}(x)$ $=$ $(-4)^{n}n!x\,_{1}F_{1}\left(\frac{1}{2}-n;\frac{3}{2};x^{2}\right),\quad n\geq 0.$
@@
The results in [19] cannot be taken as a general rule in terms of the relation between the Darboux process going from one differential operator to another and the corresponding monodromy groups. The peculiar situation in [19] is that a sequence of Darboux transformations preserves the monodromy group in two different instances, in one case the groups are all trivial and in the other case they coincide with the integers. The paper [34] looks at an example of the hypergeometric equation of Euler, Gauss and Kummer with a finite monodromy group (all of these were characterized by H. Schwartz) and one sees that an application of the Darboux process to it gives a differential operator with an infinite monodromy group.
@@
This is much simpler than the situation in the Hermite case, there is no need to do any scaling, as in the expressions for $\psi_{2n+1}$ and $\psi_{2n}$ above starting from the most familiar solutions of the differential equation, namely $P_{n}$ and $Q_{n}$.
@@ 4. Example 1: Hermite with one gap
$\mathcal{H}_{x}+2I=\partial_{x}^{2}-2x\partial_{x}+2=(\dfrac{1}{x}\partial_{x}-2)(x\partial_{x}-1)$
@@
$p_{n}(x)=\det\begin{pmatrix}x&1\\ H_{n}(x)&H^{\prime}_{n}(x)\end{pmatrix}$.
@@
$\dfrac{n(n-1)}{2}p_{n-2}(x)+\dfrac{2n-1}{4}p_{n}(x)+\dfrac{n-1}{8(n+1)}p_{n+2}(x)=\dfrac{x^{2}}{2}p_{n}(x),\quad n\geq 1,$
@@
For a small sample of references on this topic see [20, 21, 22, 27, 28, 29, 30, 49, 53, 56]. It may be worth remarking that the first use of the Darboux process in the context of exceptional orthogonal polynomials appears to be [53] where the author very graciously says “This procedure whose idea goes back to an older paper [19]”, while the first observation of the presence of recursion relations appears to be [56].
@@
highlighting the important role of the operator $x\partial_{x}-1$ that appears in the factorization of $\mathcal{H}_{x}$ above, in giving new eigenfunctions in terms of old ones.
@@
$\phi_{6}(x)$ $=$ $-120\widetilde{\alpha}_{0}\left(1+6x^{2}-12x^{4}+\dfrac{8}{3}x^{6}\right)+480\widetilde{\beta}_{2}\left(x^{3}-\dfrac{3}{5}x^{5}+\dfrac{3}{70}x^{7}+\ldots\right).$
@@ 5. Example 2: Hermite with two gaps
$q_{n}(x)=\det\begin{pmatrix}4x^{2}-2&8x\\ H_{n-1}(x)&H^{\prime}_{n-1}(x)\end{pmatrix}$.
@@
$\phi_{1}(x)$ $=$ $\widetilde{\alpha}_{1}\left(1+3x^{2}-\dfrac{x^{4}}{6}-\dfrac{13}{30}x^{6}-\dfrac{11}{56}x^{8}+\ldots\right)+\widetilde{\beta_{1}}x,$
@@ 6. The ad-conditions
$\left(\textrm{ad}\ X\right)^{n}(Y)=\textrm{ad}\ X\left(\left(\textrm{ad}\ X\right)^{n-1}(Y)\right),\quad n\geq 2.$
@@
$\textrm{ad}\ \mathcal{L}^{2}(\Theta)-4\textrm{ad}\ \mathcal{L}^{0}(\Theta)=0,$
@@
$\textrm{ad}\ \mathcal{L}^{3}(\Theta)-16\textrm{ad}\ \mathcal{L}^{1}(\Theta)=0.$